UT 20131216 LUCI2 Nighttime Commissioning

Observer: JHeidt, APramskiy, DThompson
Telescope Operator: GBechetti
LBTO Support: TSargent, DMiller

Plan for the night

- Verify AGw-transform
- LS-acquisition
- MOS-acquisition

Summary

- Longslit acquisition works

- MOS-errors three times, but could be recovered

Preparations

Ingo revised his LS/MOS-acquisition tool, did closed dome tests in the afternoon
=> offsets are now in detxy, looks promising !!!

Details

Times in UT, files are luci2.20131215.iiii

0001 0010 Some testing by Alexander

### Readnoise.xml
0011 0040 1x2.7355s LIR/NORM
0041 0070 1x6.8515s MER/NORM (@0062..0069 slew tel to horizon)

### Darks for twilight flats
0071 0075 20x2.75s LIR/INT (also test RN scaling with Ndit)
0076 0125 10x10.0s LIR/INT

### Twilight flats
0126 0135 1x2.75s FeII 23k ADU
0136 0145 1x2.75s P_beta 28k
0146 0155 1x2.75s FeII 16k
00:20 Sunset
0156 0165 1x2.75s P_beta 17k
0166 0175 1x2.75s FeII 5k
0176 0185 1x2.75s P_beta 6k

0186 0205 1x2.75s K 7...4.4k
0206 0215 1x2.75s K 4k

Weather at sunset: good, mostly clear, a few cirrcus in SW, T=4deg, hum=25%, wind 10m/s

00:58 Luci2-cooling alarm on both pressures sensors show a
spike atthe same time, both going down than

01:00 Geno did collimation and pointing check

MOS-acq. - test data for Ingo

01:05 Jaron's M225UDS3H1.xml, didn't work, need to check what wrong

01:20 M225UD3H1, MOS-acq, N1.8, preset OK but field close to moon, stopped
Geno collimates on nearby star, than we go back and take data
sky (#217), target (#218), same with N3.75, saturated, revise DIT in script

01:41 M225UD3H1, MOS-acq, N3.75, sky #220, target #221, OK
N1.8, sky #222, target #223, OK, throughmask N1.8, #224, N3.75, #225

*There seems to be an instability at the readout electonics, from time to time
there seems to be step in amplifiers, need to check the amount

LS-acqusition - single star

02:03 Preset to BS9107, PA=0, N3.75, target #226, source=1105,1121
requested=973,900, verification #227, found=983,895, OK
Redo N1.8, target #228, source=1005,966, requested=977,917, verification #229, found=982,913, OK

02:24 Preset to BS9107, PA=225, N3.75, target #230, source=1092,747
requested=1042,918, verification #231, found=1033,918, OK
Redo N1.8, target #232, source=1029,976, requested=1029,919
verification #233, found=1031,917, OK
Redo N3.75 with sky-subtraction: sky #234, target #235, source(#235-#234)=693,831
requested=855,864, verification #236, found=855,869,OK

works !!!

02:40 Seeing very variable, 0.7...2.5"

LS-acqusition - two stars

02:49 Preset to BS9107, PA=0, N3.75, #237, verification #238,
doesn't seem to work, redo, system hangs, killall -9 java, startup again, target #239, sources=1112/1124(1), 856,1398(2), requested=987,1013(1),982,1387(2)
Offset: -170pix,-26pix,-42.02deg
slitview #240, verification #241,found=930,1112(1), 559,1142(2)
=> Rotation goes in wrong direction (should be +42deg)
=> everything in Ingos SW, where rotation is involved should change sign

04:15 Alexander changes the sign in output. Check if that helps.
Seeing stabilizes below 1", some cirrus around
target #246, sources=933,1111(1),742,1403(2), requested=990,987(1),983,1246(2),
angle+42deg this time, verification#247, found=949,837(1),1009,1184(2)
=> nope, overshooting (angle too large) => must be in matrix
lets wait until Ingo is in his office

MOS-acq. - see how it goes

#248 crap, did some collimation on a bright star first

04:53 Back to M225UDS3H1, N3.75, sky #249, target #250, slitview #251
Alignment and verification #252, did something wrong
Did some tests and found the problem (one of the slitlets on top), redo MOS-acq.

05:52 MOS-error when trying to move from FPU to turnout, holding card 1 motor 5 MHU head translation
Need to contact Peter

06:20 More cirrus coming in, sky pretty much covered, temp rose up to +8deg, humidity = 5%

07:00 Peter has put the MOS back to live

LS-acqusition - two stars

07:14 Alexander deployed a new version of Ingos tool, check if LS-acq with 2 stars works now

07:24 Preset to BS9126, PA=45, test for LS-acq. 2 stars, no two stars to center, move the field a bit,
tried to to alignment, output far wrong (offset degrees), screen hangs, restart SW, new image
of field #257, tried to align, results again crappy (but used LS-image taken at another region on the sky)
do new LS-image (#258), image field => GEIRS got stuck, killall -9 java, set up everything new,
tried various times to align two stars with the slit, no success .

09:55 Got another MOS-error, this time when moving 2"-slit from storage to turnout

10:45 Recovered by Hans & Walter

10:52 MOS-error again, recovered

11:15 Ingo provided a further version of his acquisition tool

Did successfully a variety of tests, with aligning one/two sources in the slit smile
Once a successfull AGw-transform is in place it should work perfectly.
At present for each preset/position angle a separate throughslit image is required
when 2 stars shall be aligned. Thsi must be fixed. Details:
BS9126, #270 source image, #271 blind image, #272 slit ima, offset applied
#273 throughslit, no source, #274 mask in turnout, #275 OK, after offset of x = 1.5"
BS9139, #276 slit ima, #277 source ima, offset -4.5",-3.1", #278 verification,
#279 through slit, src in slit, #280/281 further throughslist
BS9139, different PA, #282 slit ima, #283 src ima, resent to PA=55,
#284 src ima, #285 slit ima , #286 applied offset, #286 src not in slit, #287 src ima for delta x offset (small), #288 src in slit

works !!!

13:18 BS9142, test Michelle's script, LS-alignment for throughput-determination
#289 sky, #290 src, #291 slitview, #292 alignment good, #293-304 taking
test spectra in 210zJHK, stopped when guider stopped WFS

14:55 Closing, cirrus outside, no flats

-- JochenHeidt - 16 Dec 2013
Topic revision: r8 - 16 Dec 2013, JochenHeidt
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback